sanjay is now 1/2 as old as his brother. In 6 more years he will be 3/5 as old as his brother. Find the present age of each boy.
Answer by RAY100(1637)

Let s = Sanjay age,,,b= Brother's age
.
(1) s = (1/2) b
.
(2) ( s+6) = (3/5) (b+6)
.
subst (1) in (2)
.
(1/2) b +6 = (3/5)b + 18/5
.
6- 18/5 = (3/5 -1/2) b
.
2.4 = .1 b
.
24 = b,,,,,,s= 1/2 b = 1/2 *24 = 12
.
check
(1) 12 = 1/2 * 24 ,,,,ok
.
)2) 12 +6 = (3/5) ( 24 + 6 )
.
18 = 3/5 *30 =18 ,,,,,,ok